OpenStudy (anonymous):

Write an equation of a line that passes through the point (8, 4) and is parallel to the line y = 4x + 2.

OpenStudy (anonymous):

I cant remember the formula

OpenStudy (anonymous):

y-y0 = m(x-x0)

OpenStudy (anonymous):

point (8, 4): x0=8, y0=4

OpenStudy (anonymous):

y-4=m(x-8)

OpenStudy (anonymous):

yep :) you also know m, because parallel lines......

OpenStudy (anonymous):

m=2

OpenStudy (anonymous):

y = 4x + 2. y = mx + c

OpenStudy (anonymous):

4 lol my bad

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

right and then re/arrange and get the second line in standard form

OpenStudy (anonymous):

y-4=4x-28

OpenStudy (anonymous):

4x8 = 32

OpenStudy (anonymous):

messed up lol answer is y=4x-28

OpenStudy (anonymous):

ahhh

OpenStudy (anonymous):

right

OpenStudy (anonymous):

wrong direction though ?

OpenStudy (anonymous):

thank you so much hun!! (:

OpenStudy (anonymous):

ah no its right ;)

OpenStudy (anonymous):

yw